### v2

# GloTok: Global Perspective Tokenizer for Image Reconstruction and Generation

Abstract
Existing state-of-the-art image tokenization methods leverage diverse semantic features from pre-trained vision models for additional supervision, to expand the distribution of latent representations and thereby improve the quality of image reconstruction and generation. These methods employ a locally supervised approach for semantic supervision, which limits the uniformity of semantic distribution. However, VA-VAE proves that a more uniform feature distribution yields better generation performance. In this work, we introduce a Global Perspective Tokenizer (GloTok), which utilizes global relational information to model a more uniform semantic distribution of tokenized features. Specifically, a codebook-wise histogram relation learning method is proposed to transfer the semantics, which are modeled by pre-trained models on the entire dataset, to the semantic codebook. Then, we design a residual learning module that recovers the fine-grained details to minimize the reconstruction error caused by quantization. Through the above design, GloTok delivers more uniformly distributed semantic latent representations, which facilitates the training of autoregressive (AR) models for generating high-quality images without requiring direct access to pre-trained models during the training process. Experiments on the standard ImageNet-1k benchmark clearly show that our proposed method achieves state-of-the-art reconstruction performance and generation quality.
1 Introduction
Recent years have witnessed remarkable advancements in image generation, driven by autoregressive (AR) models (DBLP:journals/corr/VaswaniSPUJGKP17; radford2018improving; touvron2023llama) and diffusion models (dhariwal2021diffusion; ho2022classifier; rombach2022high). To promote high-quality generation, many contemporary methods adopt a two-stage paradigm: (1) a tokenizer compresses input images into tokens, followed by (2) a generator modeling the distribution of tokens. The compressed tokens determine both structural composition and detailed characteristics of generated images. Therefore, image tokenizers play a crucial role in improving the performance of image generation.
Figure 1: Comparison of generative performance among different methods and GloTok, where lower values on the Y-axis correspond to better performance.
Recent SOTA tokenizers (zhu2024stabilize; qu2024tokenflow; li2024imagefolder; bai2024factorized; ma2025unitok) for image generation with AR models leverage pre-trained models such as DINO (DBLP:journals/corr/abs-2104-14294; oquab2023dinov2; darcet2023vitneedreg) and CLIP (radford2021learning) as additional supervision. Concretely, the pre-trained model converts the input images into semantic features, and these features are then served to guide the tokenizer to generate more effective latent representations for image generation.
While these approaches have achieved remarkable performance in enhancing the information capacity of latent representations, their dependence on single-image semantic supervision poses a significant challenge to semantic codebook fitting. Specifically, this local approach is confined to performing contrastive learning between latent-space features and pre-trained features within individual images, thereby constraining its capacity to model feature discrepancies across the entire dataset. Consequently, the learned latent-space distribution exhibits suboptimal breadth and uniformity. VA-VAE (yao2025reconstruction) demonstrates that more uniform latent representations can significantly improve generative performance. Based on the conclusions of this prior work, suboptimal uniform feature distributions can tend to constrain generative performance.
In this paper, we introduce GloTok, a novel tokenizer that outperforms previous work in Fig. 1 and is designed to achieve a more uniform latent distribution, leveraging global relations across the entire dataset. In practice, GloTok adopts a dual codebook architecture to learn the semantics and details of the input images. A codebook-wise histogram relation learning method is proposed to transfer the semantics to the semantic codebook by feature relationships modeled by pre-trained models across the entire dataset (called global relations). Taking ImageNet as an example, existing methods require batch-based semantic learning for 1.28 million images, resulting in unstable optimization and makes it difficult to achieve global optimality. Differently, our approach consistently leverages global relations as the semantic guidance, such as a multi-bin histogram to summarize the global relations. This design not only facilitates improved uniformity of semantic latent distribution, but also eliminates the need to directly access the pre-trained model during training, thereby reducing both temporal and spatial overhead.
In addition, to mitigate accuracy degradation caused by discretization, we employ a residual learning module in GloTok. This module takes quantized features as input, predicts the residual between the original continuous features and their quantized discrete counterparts, and fuses this residual with the quantized features to produce the final quantized feature representations. Residual learning enables the tokenizer to partially restore fine-grained details, thereby improving the quality of image reconstruction.
Through a novel histogram relation learning method and residual learning modules, the performance of the AR model trained with our proposed GloTok achieves significant improvements in generative tasks. Our experiments demonstrate that the proposed model achieves a state-of-the-art (SOTA) reconstruction FID of 0.83 at $\times256$ resolution on the standard ImageNet benchmark. And we get a competitive generation FID=1.75 on the ImageNet $\times256$ image generation using an autoregressive generator.
Our contribution can be summarized as follows.
  • We propose GloTok, a novel tokenizer that constrains the global relational distribution within the latent space to transfer the semantics from pre-trained models, achieving a significant improvement in the performance of image reconstruction and generation.
  • A codebook-wise histogram relation learning method is introduced to transfer the semantics to the semantic codebook, which models a more uniform latent distribution and eliminates the need to directly access the pre-trained models during training.
  • To mitigate accuracy degradation induced by discretization, a residual learning module is designed to predict and fuse token residuals into the final representations, preserving fine-grained details for superior image quality.
2 Related Work
2.1 Image Tokenization
Image tokenization is the fundamental process in image reconstruction and generation. An image tokenizer converts the input image into continuous or discrete tokens.
As a pioneering work on transformer-based image generation, VQ-GAN (esser2021taming) proposes the encoder-quantizer-decoder architecture, which employs a vector quantization (VQ) technique to map the continuous latent space features of the input images into discrete tokens. However, the performance of VQGAN is constrained by both a small codebook size and a low token utilization rate. ViT-VQGAN (yu2021vector) and Efficient-VQGAN (cao2023efficient) leverage the Vision Transformer (ViT) architecture (dosovitskiy2020image), which mitigates reliance on convolutional inductive biases due to its global attention mechanisms. RQ-VAE (lee2022autoregressive) adopts a residual quantization (RQ) to precisely approximate the feature map. VQGAN-LC (zhu2024scaling) and SimVQ (zhu2024addressing) employ linear layers to expand the codebook capacity while preserving high codebook utilization efficiency. Furthermore, MAGViT-v2 (yu2023language) proposes a lookup-free quantization method leading to a large vocabulary. Unlike typical 2D tokenizers, TiTok (yu2024image) adopts a 1D tokenization framework that enables a more flexible token count design while also capturing semantically rich image information.
Recent advancements have focused on injecting semantic features from pre-trained vision models such as DINO (DBLP:journals/corr/abs-2104-14294; oquab2023dinov2; darcet2023vitneedreg) or CLIP (radford2021learning) into the codebook to enhance the generation performance of VQGAN. DiGiT (zhu2024stabilize) integrates the discrete tokens derived from clustered DINO (oquab2023dinov2) features into the latent space of VQGAN to enrich the semantic information of the latent space of the generative model. UniTok (ma2025unitok) adopts a pre-trained CLIP (radford2021learning) model to align the latent space of input images with the semantic latent space derived from the corresponding text. Recent work, such as ImageFolder (li2024imagefolder) and FQGAN (bai2024factorized), utilizes a multi-codebook architecture in which certain codebooks are designed to capture the feature distributions extracted from pre-trained vision models using a feature-wise contrastive loss. These methods adopt a local perspective, specifically by capturing semantics through intra-image feature comparison with pre-trained models. In contrast, our tokenizer GloTok employs a method that globally constrains the distribution of semantic features with pre-trained relational distributions to transfer the semantics.
2.2 AR Image Generation
Recent years have witnessed remarkable advancements in auto-regressive (AR) image generation, where models synthesize images by capturing the distribution of discrete tokens. Early foundational works, such as VQGAN (esser2021taming), VIT-VQGAN (yu2021vector), and RQ-Transformer (lee2022autoregressive), employ transformer architecture to model discrete token sequences for image generation. Building on this, LlamaGen (sun2024autoregressive) introduces the Llama (touvron2023llama) model to enhance image generation capabilities. Beyond architectural innovations, optimizing token training strategies has emerged as a key research direction. For instance, VAR (tian2024visual) adopts a multi-resolution token prediction framework to improve the image fidelity and detail preservation. MAR (li2024autoregressiveimagegenerationvector) models the per-token probability distribution using a diffusion procedure, which achieves strong generation results. FlowAR (ren2024flowar) introduces a general next-scale prediction method with a streamlined scale design. RAR (yu2024randomized) randomly permutes the input token sequence into different factorization orders for modeling bidirectional contexts. xAR (ren2025nexttokennextxpredictionautoregressive) enables flexible prediction units and effectively alleviates exposure bias, achieving significant generative performance.
3 Method
3.1 Preliminary
AR image generation models are trained from discrete tokens generated from an image tokenizer. The tokenizer typically follows an encoder-quantizer-decoder architecture, where the encoder extracts features from the image, the quantizer converts these features into discrete tokens, and the decoder reconstructs the image from these tokens. This process allows the model to effectively learn and generate images based on tokenized representations.
Given an image $x\in\mathcal{R}^{H\times W\times 3}$, the encoder $E$ first extracts continuous features $Z^{h\times w\times c}$ from the input image $x$, where $h\times w$ represents the sequence length of the latent tokens and c represents the channel dimension. For image quantization, the quantizer selects the closest token $t$ for each feature $z$ of $Z$ with a learnable codebook $C^{n\times d}$ by calculating the Euclidean distance (esser2021taming) or the cosine similarity (shi2024taming; unitok; bai2024factorized). The quantization loss $L_{Q}$ is calculated based on the quantized features $\hat{Z}^{h\times w\times c}$ and the original features $Z$ to optimize the latent space and the codebook:
$L_{Q}=\left\|sg\left[Z\right]-\hat{Z}\right\|_{2}^{2}+\beta\left\|sg\left[\hat{Z}\right]-Z\right\|_{2}^{2}$ (1)
Here, sg[·] denotes the stop-gradient operation and $\|\cdot\|_{2}^{2}$ denotes the Euclidean distance (or called L2 distance). The quantization process optimizes latent representations within the codebook while reducing the representation gap between the continuous latent space and the discrete latent space.
After quantization, the decoder $D$ decodes $\hat{Z}$ into the reconstruction image $\hat{x}$. For the representative VQGAN (esser2021taming) model, the optimization of the entire network is represented by the following loss function $L_{VQGAN}$:
$L_{R}=\left\|x-\hat{x}\right\|^{2}$ (2)
$L_{VQGAN}=\lambda_{R}L_{R}+\lambda_{Q}L_{Q}+\lambda_{P}L_{P}+\lambda_{G}L_{G}$ (3)
where $\lambda_{i}$ is the weight factor for each loss term, $L_{R}$ is the reconstruction loss, $L_{P}$ is the perceptual loss between $x$ and $\hat{x}$, and $L_{G}$ is adversarial loss encouraging to learn more realistic images $x$ with a patch-discriminator  (isola2017image).
Figure 2: Illustration of our method. Top: GloTok encoder-quantizer-decoder architecture with dual codebooks and residual modules. Bottom: overview of the Histogram Relation Learning method. An image is quantized into two sets of tokens by a visual codebook and a semantic codebook. The semantic codebook learns the token relationship from features clustered from a pre-trained model with a histogram loss. GloTok adopts two residual modules to learn the residuals between continuous features and discrete features.
3.2 Architecture
Similarly to previous methods (bai2024factorized; li2024imagefolder), as shown in Fig. 2, we leverage a dual-codebook architecture with a convolutional encoder and decoder. we encode the input image into two sets of tokens: one consisting of the semantic tokens and the other of the visual tokens. We use the semantic tokens to model the distribution of semantic features, while visual tokens are employed to model the image’s visual features. Specifically, the encoder $E$ first encodes the input image $x$ into the latent features $Z^{h\times w\times c}$. We then adopt two adapters to decompose $Z$ into semantic features $Z_{sem}$ and visual features $Z_{vis}$. For quantization, $Z_{sem}$ and $Z_{vis}$ are quantized into $\hat{Z}_{sem}$ and $\hat{Z}_{vis}$ respectively, with learnable codebooks $C_{sem}$ and $C_{vis}$.
To control the global feature space distribution of semantic features, we introduce a novel codebook-wise histogram relation learning method that constrains the distribution of relations between tokens in the semantic codebook, detailed in Section 3.3. The visual codebook is specifically served to capture visual feature representations within the latent space, enabling image reconstruction by leveraging structural patterns learned from the semantic codebook.
Quantization of the continuous latent $Z$ into discrete representations inherently induces information degradation. To mitigate the limitation, we designed a residual learning module to recover lost information, thus improving the quality of image reconstruction. The learned residual features are aggregated with quantized tokens through summation operations to generate information-complete token representations. Subsequently, the decoder decodes the composite latent representation $\hat{Z}$ formed by channel-wise concatenation of $\hat{Z}_{sem}$ and $\hat{Z}_{vis}$ to synthesize the output image.
3.3 GloTok
Histogram relation learning
In order to visualize the distribution of the quantitative data, the histogram is a commonly used method that reflects the probability density of the underlying data distribution. Inspired by this concept, we propose a codebook-wise histogram relation learning method that transfers the prior global relational distribution knowledge to the semantic codebook.
To achieve this objective, it is necessary to select an appropriate source for the prior distribution. We adopt two approaches for its construction: (1) directly utilizing the semantic codebook of the tokenizer from a pre-trained model; (2) performing k-means clustering on the intermediate layer features of CLIP (radford2021learning) or DINOv2 (oquab2023dinov2) adopted in VQGAN-LC (zhu2024scaling).
In particular, our methodology differs from VQGAN-LC in that it employs an MLP layer to establish the mapping from clustered features to the codebook. Although the feature sources in VQGAN-LC contain rich semantic information, the semantic constraints imposed on the mapped codebook are insufficient, and the use of an MLP layer for training a large codebook, which is up to 100,000, results in substantial computational cost. Instead, we focus on the patterns of relational similarity between tokens.
To construct the global relational distribution, we first compute the pairwise cosine similarities between tokens within each token set $C^{n\times d}$:
$D=\{\left.f(t_{i},t_{j})\right|t_{i},t_{j}\in C,i\leq j\}$ (4)
where f[$\cdot$] is a cosine similarity function. Subsequently, in a histogram-like manner, we initialize N bins slicing from -1 to 1, where each bin corresponds to a specific sub-range of quantized token distances:
$B=\left\{\left.-1+\frac{2k}{N-1}\right|k=0,1,2,...,N-1\right\}$ (5)
To quantify the distribution of $D$ across different intervals $B$, we compute a Gaussian-smoothed histogram of pairwise cosine similarities:
$p_{n}=\sum_{i=1}^{\|D\|}\exp{(-\alpha(D_{i}-B_{n})^{2})}$ (6)
where $\|D\|$ is the dim of D and $\alpha$ is a weight factor.
We define the relational distribution of the trainable codebook as the student distribution $P_{s}$ and define the relational distribution of pre-trained tokens as the teacher distribution $P_{t}$. Then, we employ a Kullback-Leibler divergence to constrain the student distribution with the teacher distribution:
$L_{hist}=KL(\log{P_{s}}\|sg\left[P_{t}\right])$ (7)
where sg[·] denotes the stop-gradient operation, $P_{s}$ denotes the normalized student distribution, and $P_{t}$ denotes the normalized teacher distribution.
Residual learning
The quantization process, which transforms continuous features into discrete representations, inherently induces precision degradation. While discrete representations simplify the latent space expression, they inevitably compromise the model’s representational capacity.
We propose a residual learning method in which residual modules $R_{sem}$ and $R_{vis}$ are implemented as transformer blocks process quantized latent features $\hat{Z}_{sem}$ and $\hat{Z}_{vis}$, respectively, to predict lost latent details during discretization:
$Z_{res,sem}=R_{sem}(\hat{Z}_{sem})$ (8)
$Z_{res,vis}=R_{vis}(\hat{Z}_{vis})$ (9)
The optimization is derived from the differential mapping between continuous features $Z$ and the quantized features $\hat{Z}$ through residual error minimization:
$\begin{split}L_{res}&=\left\|sg\left[Z_{sem}\right]-\hat{Z}_{sem}-Z_{res,sem}\right\|\\ &+\left\|sg\left[Z_{vis}\right]-\hat{Z}_{vis}-Z_{res,vis}\right\|\end{split}$ (10)
[…]
$\hat{Z}_{vis}=\hat{Z}_{vis}+Z_{res,vis}$ (11)
$\hat{Z}_{sem}=\hat{Z}_{sem}+Z_{res,sem}$ (12)
Training loss
Our model is trained with a combined loss:
$L=L_{VQGAN}+\lambda_{res}L_{res}+\lambda_{hist}L_{hist}$ (13)
where $\lambda_{res}$ and $\lambda_{hist}$ are weight factor that respectively control $L_{res}$ and $L_{hist}$, thus balancing various loss terms.
3.4 Training with Auto-Regressive Model
Through quantization, the input image is discretized into two complementary token sets in our tokenizer. Previous auto-regressive models (esser2021taming; touvron2023llama; tian2024visual) are trained with the indices of the tokens quantized by codebook. These auto-regressive models are designed to model inter-token dependencies via sequential pattern learning.
However, although previous multi-branch models (bai2024factorized; li2024imagefolder) exhibited superior reconstruction performance compared to other single-branch methods (sun2024autoregressive; tian2024visual), their generation performance remained relatively suboptimal. When the tokenizer employs a multi-branch architecture, it must generate multiple indices to represent images. Due to the exposure bias (arora2023exposurebiasmattersimitation) inherent in AR training, producing multiplicative indices exacerbates index prediction errors during inference, thus degrading generation performance. In our work, we aim to circumvent this limitation.
We concatenate the semantic features and visual features obtained after vector quantization along the channel dimension, and employ an xAR (ren2025nexttokennextxpredictionautoregressive) model to directly learn this combined feature space. This approach enhances the generation performance of the multi-branch tokenizer.
Codebook Type Method Codebook Size rFID $\downarrow$ PSNR $\uparrow$
Single VQGAN (esser2021taming) 16,384 5.96 23.3
Single RQ-VAE (lee2022autoregressive) 16,384 3.20 -
Single VQGAN-LC (zhu2024scaling) 16,384 3.01 23.2
Single VQGAN-LC (zhu2024scaling) 100,000 2.62 23.8
Single LlamaGen (sun2024autoregressive) 16,384 2.19 20.7
Single MaskBit (weber2024maskbit) 16,384 1.37 21.5
Signle TiTok-S-128 (yu2024image) 4,096 1.71 17.7
Single Open-MAGVIT2 (luo2024open) 16,384 1.58 -
Single Open-MAGVIT2 (luo2024open) 262,144 1.17 22.6
Single IBQ (shi2024taming) 16,384 1.37 21.0
Single IBQ (shi2024taming) 262,144 1.00 -
Muti TokenFlow (qu2024tokenflow) $32,768\times 2$ 1.37 21.4
Muti FQGAN-Dual (bai2024factorized) $16,384\times 2$ 0.94 22.0
Muti GloTok $16,384$ 0.92 22.4
Muti GloTok(2x) $16,384\times 2$ 0.83 22.7
Table 1: Reconstruction comparison with other tokenizers evaluated on ImageNet 256×256 50k validation dataset.
4 Experiments
4.1 Experiments Setting
Implementation details
We trained and evaluated GloTok on the $256\times 256$ ImageNet (deng2009imagenet) benchmark. The encoder and decoder settings are consistent with VQGAN (esser2021taming). To facilitate alignment in methods employing single-branch codebook and multi-branch codebook settings with the same codebook size, respectively, we trained two tokenizers with distinct codebook size configurations: (1) the semantic codebook is configured with a size of 4,096, and the visual codebook with a size of 12,288, resulting in a total size of 16,384; (2) both the semantic and visual codebooks have a size of 16,384, which we refer as GloTok(2x). The dimension of the codebooks is 8. To transfer inter-token relational knowledge from pre-trained models, we train the tokenizer with diverse teacher distributions, such as K-means-clustered DINOv2 tokens and a pre-trained semantic codebook. We configure the number of bins (B) of the histogram learning method to 40. We set $\lambda_{hist}=0.01$, $\lambda_{res}=0.5$. Within the residual learning module, we incorporate a single-layer transformer block to predict residual components. The tokenizer is trained with the following settings: a learning rate of 2e-4 without any decay mechanism, an Adam Optimizer with $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.95, a global batch size of 240 with 150 epochs. For the generator, we train FAR-B, FAR-L with 200 epochs with our tokenizer, and we train xAR-B, xAR-H  (ren2025nexttokennextxpredictionautoregressive) with 800 epochs following xAR settings.
Metrics
We employ the Frechet Inception Distance (FID) (DBLP:journals/corr/HeuselRUNKH17) and the Peak Signal-to-Noise Ratio (PSNR) to evaluate reconstruction quality. We use the ImageNet validation set, consisting of 50k samples, to compute the reconstruction FID (rFID) and PSNR. To evaluate the quality of generated samples, we employ two widely recognized metrics: generation FID (gFID) and Inception Score (IS) (salimans2016improved). The gFID measures the similarity between the distribution of generated images and that of real-world data, providing a quantitative assessment of the realism and diversity of the outputs. On the other hand, the IS evaluates both the quality and diversity of the generated samples by leveraging an Inception model pretrained on the ImageNet dataset.
[…]
AR VQGAN (esser2021taming) 227M 256 18.65 80.4
AR VQGAN-LC (zhu2024scaling) 404M 256 15.4 -
AR LlamaGen-B (sun2024autoregressive)  [^1] M 256 5.46 193.61
AR LlamaGen-XXL (sun2024autoregressive) 1.4B 256 3.09 253.60
AR LlamaGen-3B (sun2024autoregressive) 3.1B 256 3.06 279.71
AR IBQ-B (shi2024taming) 342M 256 2.88 254.73
AR IBQ-XXL (shi2024taming) 2.1B 256 2.05 286.73
VAR ImageFolder (li2024imagefolder) 362M 256 2.60 295.0
VAR VAR-d16 (tian2024visual) 310M 256 3.30 274.4
VAR VAR-d24 (tian2024visual) 1.0B 256 2.09 312.9
VAR VAR-d30 (tian2024visual) 2.0B 256 1.92 323.1
FAR FQGAN (bai2024factorized) 415M 512 3.38 248.26
FAR FQGAN (bai2024factorized) 898M 512 3.08 272.52
FAR GloTok 415M 512 3.13 280.66
FAR GloTok(2x) 415M 512 2.95 261.38
FAR GloTok 898M 512 2.98 285.34
xAR VQ-VAE† 172M 256 2.47 277.63
xAR GloTok 172M 256 2.12 272.16
xAR GloTok(2x) 172M 256 2.00 269.25
xAR GloTok 1.1B 256 1.75 327.50
Table 2: Class-conditional generation comparison on ImageNet-1k 256×256 resolution. The evaluation protocol and implementation are the same as ADM. †denotes that the xAR is trained with a VQ-VAE tokenizer from LlamaGen (touvron2023llama).
4.2 Main Results
Image reconstruction
As shown in Tab. 1, our tokenizer was compared with other state-of-the-art models in terms of reconstruction performance. Reconstruction performance of our tokenizer was assessed at $\times256$ resolution, employing 16x downsampling, on the ImageNet-1K validation set.
[…]
Figure 3: ImageNet-1k $\times256$ generated samples of GloTok trained with xAR.
Image generation
[…]
4.3 Histogram Relation Learning Makes More Uniform
[…]
VQ-VAE† 0.221 0.9971 0.1191
FQGAN 0.225 0.9970 0.1202
GloTok 0.217 0.9972 0.1168
Table 3: Comparison of uniformity metrics of VQ-VAE, FQGAN, GloTok. †denotes that the VQ-VAE is obtained from LlamaGen (touvron2023llama).
4.4 Ablation Studies
To validate the effectiveness of our proposed framework, we conducted ablation studies on GloTok with a 4096 semantic codebook size and a 12288 visual codebook size.
The impact of different feature levels
As illustrated in Fig. 4, to evaluate the impact of different feature levels on overall image reconstruction, we visualize how varying retention ratios of visual features (ranging from partial to full preservation) influence image quality both with and without semantic feature guidance. Specifically, we set the discarded features to 0. We retain 100% of the semantic features while varying the retention ratio of the visual features from 25% to 100% to observe their interaction.
The results show that visual features alone preserve substantial color information in reconstructed images, while structural and textural details are diminished. However, when semantic information is incorporated, the structural and textural components are effectively restored. The difference indicates that the semantic codebook integrates high-frequency components (e.g., structural and textural elements), while the visual codebook combines low-frequency components (e.g., color distributions). Our experiments demonstrate the effectiveness of the global relational learning method in transferring semantics into the codebook.
The effect of the histogram relation learning
Our tokenizer is trained for 40 epochs under varying weight configurations of histogram relation learning to investigate its impact on the reconstruction performance of our tokenizer. As shown in Tab. 4, our experimental analysis reveals that incorporating the method facilitates a reduction in rFID from 1.43 to 1.21, accompanied by an improvement in PSNR from 22.06 to 22.23. These findings demonstrate that histogram relation learning plays a pivotal role in enhancing tokenizer reconstruction performance. Empirically, setting the hyperparameter weight to 0.01 achieved superior optimization outcomes in our experiments.
Figure 4: Visualization of the reconstruction on different retention rates of visual features. Each row presents visualization results for visual features with different retention ratios, both with and without semantic features.
[…]
$\times$ - 1.43 22.06
$\checkmark$ 0.1 1.27 21.97
$\checkmark$ 0.01 1.21 22.23
Table 4: Effect of the histogram relation learning. Tokenizers are only trained by the histogram relation learning method without the residual module.
Different components
We evaluated the contributions of individual components to the improvement of GloTok reconstruction. Results are reported in Tab. 5. We start with a vanilla dual-codebook design. Through the adoption of histogram relation learning, our experiments demonstrate a significant improvement in reconstruction quality, with the Fréchet Inception Distance (FID) decreasing from 1.43 to 0.97 and the Peak Signal-to-Noise Ratio (PSNR) increasing from 22.06 to 22.30. These findings show that histogram relation learning enhances the diversity of the latent space. We then incorporate the residual learning module, which improves the rFID from 0.97 to 0.92 and improves the PSNR from 22.30 to 22.44. These results demonstrate the effectiveness of the residual learning module in optimizing the reconstruction performance of the tokenizer.
Histogram Relation Residual rFID $\downarrow$ PSNR $\uparrow$
$\times$ $\times$ 1.43 22.06
$\checkmark$ $\times$ 0.97 22.30
$\checkmark$ $\checkmark$ 0.92 22.44
Table 5: Ablation study on different components for the improvement of GloTok.
5 Conclusion
In this paper, we introduce GloTok, an image tokenizer that leverages global relational information to achieve a more uniform semantic latent distribution. Employing a codebook-wise histogram relation learning method, GloTok effectively regulates the global distribution of tokens across the entire dataset. GloTok does not have to directly access the pre-trained model during training, effectively reducing both training time and GPU memory usage. In addition, a residual learning module is proposed to mitigate the fine-grained detail lost during discretization. Experiments on the ImageNet dataset show that GloTok achieves state-of-the-art reconstruction performance and generative performance.
Acknowledgments
The computations in this research involved with Yuxi Mi and Shuigeng Zhou were performed using the CFFF platform of Fudan University.